OpenStudy (anonymous):

If f(n) = n 2 - n, then f(-4) is _____. -20 20 12 -12

OpenStudy (anonymous):

A,B,C,D PLEASE HELP MEEE

OpenStudy (anonymous):

f(-4)=(-4)^2-(-4) f(-4)=16+4=20 if n^2 is in question @samtab

OpenStudy (anonymous):

replace n with -4 in f(n) = n 2 - n..

OpenStudy (anonymous):

@samtab got it now..??

OpenStudy (anonymous):

if it is n2 then -4(2)-(-4)=-8+4=-4

OpenStudy (anonymous):

KINDA IM STARTING TO just a little bit

OpenStudy (anonymous):

@Kashan its -4 if n2

OpenStudy (anonymous):

thanks for helping me :)

OpenStudy (anonymous):

you'll get...